OpenStudy (anonymous):

what is 2+x/5x-x-2/5x expressed in simplest form?

razor99 (razor99):

hello u want the steps

OpenStudy (anonymous):

yess please

OpenStudy (anonymous):

Can you use some parentheses to make that equation clearer please?

OpenStudy (anonymous):

okayy :) uhm what is (2+x)/(5x)-(x-2)/(5x) expressed in simplest form?

OpenStudy (anonymous):

Okay, thanks. So we have \[\frac{2+x}{5x}-\frac{x-2}{5x}\]Right? Since those are fractions that have the same denominator, that means we can go ahead and combine them, and we get \[ \frac{2+x-(x-2)}{5x} \]Can you simplify it from there?

OpenStudy (anonymous):

uhm lemme try

OpenStudy (anonymous):

nvm i got it! lol thanks!